Consider the titration of a 25.0mL sample of 0.125M HCN (Ka = 6.17x10^-10) with 0.160 M KOH.

1. What is the initial pH?

2. What is the pH at 6.00 mL base added?

3. What is the pH when 19.53 mL of KOH is added?

4. What is the PH if 21.5 mL of KOH is titrated with the solution?
Answer #1

25ML ef o. 125 M HCN 25m x o125 mel 3125 mmd HCN Ka = 6.17 xl0to Pka-log ka9 2 HCH Ht +CN t-o eqm -Ca Co Col Co2 Ka =(Callca)= 3.125mmol 0.96 m mol = 2.165mmd instrally reactcd CHCNT=2.165mpnc O.0698 M PH=pka+ log [salt]KCN TAidJ HCN + lag 9.21 O.069Snce nauid and Cid Vy less hich can gnored js be 3 1248 /PH= 9.7057 Thus,pH at 19.53 mL KoH addite is Frerd found to ka 9.7os21.5 mレ 0.16 M KoH 2l.Sm x 0.16 mol 3.44 mmol KoH Kot 3.125 mmol will heait with 3.125 mmol HuN to o 3125 Mmol KCN Thus (nKoePOkglo] = -lg 742 53 2117 ph:14-poH 83 ThpH-1.83 KoH, pH Aptai encen is fond to be 83 addilten